Marvin is trying to find the height of a tree in the school yard. He is using the proportion __Marvins's Height__ = ___tree's height_
__Marvin's shadow length_ ___tree's shadow length__ (I couldnt get it to draw a line to where it would look like it is something on top over something on the bottom)
His height is 4 feet. His shadow length is 15 inches. The length of the tree's shadow is 12 feet. Marvin used the proportion 4/15 = tree/12. But that gave the tree's height as being shorter than Marvin's. What went wrong?

Answer by rfer(16322)

you need common values.
15 inches and 12 feet doesen't work.
15 in = 1.3 ft
1.3/4=12/x
1.3x=48
x=48/1.3
x=37 ft